# arXiv:2603.01211v1[cs.AI] 01 Mar 2026

### v1

@@ 3 Operationalization: Indicators of Cultural Capabilities
Broadly, these indicators fall into two categories that balance objective verification with subjective human judgment: The first type is knowledge-based indicators, that measures the correctness and coverage of model responses against verifiable external truth sources. The goal here is to check the model’s factual recall and the fidelity of its representation of established, explicit cultural information. For example, a knowledge-based assessment might query a model on the recognized meaning of a specific national holiday, the ingredients of a regional dish, the historical significance of a landmark, detect the cultural rituals depicted in an image. Since these assessments seek answers against a known “ground truth,” they often provide an objective, scalable measure of the model’s acquired knowledge. (See Section 4 for the discussion on where scoring is a challenge even when there exist objective answers.) The second, more nuanced type is the perception-based assessment that recognizes the aspects of cultural intelligence that resides in pragmatics and context, which cannot usually be checked against a simple database. It relies on evaluating model responses using fluid, often subjective rubrics that capture perceptions of quality and fluency beyond mere factuality. These indicators move the questions from the more objective “is the answer right?” to the more subjective “is the answer appropriate?”, “is the answer fluent?” etc. This evaluation usually requires human ratings (or autoraters trained or tuned on human ratings), and are crucially dependent on the representational diversity of the human rater pool itself [25, 5, 68]. Without diverse raters, a model might be rated as “fluent” simply because it aligns with the dominant culture’s expectations, while failing to resonate with the target community.
@@
By strategically combining these knowledge-based (objective) and perception-based (subjective) indicators, researchers and practitioners can construct a robust, multi-layered “thick evaluation” [64], that moves beyond the “thin” view of mere factual accuracy or trivia questions, which may confirm a model recognizes a cultural artifact, to probe the deeper, sociocultural resonance of its outputs. This comprehensive framework ensures that the resulting measure reflects not only what a model knows, but also how effectively and safely it can act across diverse cultural contexts.
@@
Translating abstract capabilities into specific behavioral or output-based markers yields us a few key advantages. First, it offers the flexibility for developers to weigh different indicators differently depending on the overarching product goals. For instance, the global diversity indicator may be de-prioritized in the context of a model that is aimed to be launched within a specific country. Stereotype avoidance may be an indicator that is chosen based on the normative objectives of the developer. Second, it offers a reproducible methodology to assess cultural intelligence in an extensible manner. As AI models gain new modalities and as cultural norms evolve over time, our framework allows for the continuous integration of new indicators without requiring a total overhaul of the underlying capability definitions, enabling our evaluation methods to remain relevant, rigorous, and responsive over time.
@@ 4 Measurement
So far, we have defined cultural intelligence in terms of capabilities that provide the necessary conceptual clarity, as well as their corresponding indicators that offer flexible operationalization. We move now to the practical considerations of measuring these indicators at scale, a process that requires navigating the complex array of technical and sociotechnical challenges.
Consider, for instance, Factual Accuracy, an indicator of Epistemic Fidelity. Measuring this effectively requires at least three core elements: first, we must possess the ground-truth knowledge the model is being tested for; second, we need a robust a set of probes (or prompts) that necessitate the model to use or surface that fact; and finally, we require a metric that can assess how well the model’s response matches the ground truth. Each of these steps require careful consideration tailored to the specific indicator. Consequently, moving from the theoretical framework to a functional measurement requires navigating the friction between the fluid nature of culture and the relatively more rigid requirements of algorithmic benchmarking.
The challenges correspondingly coalesce around three core pillars. First, the foundational requirement of ensuring that the cultural knowledge data powering the evaluations maintains high quality, broad coverage, and a representation of necessarily pluralistic perspectives. Second, a robust evaluation design that ensures that the probes are resilient against variations in language, dialect, and structure to prevent results from being skewed by spurious linguistic patterns. And finally, the development and aggregation strategies of domain-specific metrics that can account for the unique idiosyncrasies of cultural domains, where standard NLP assumptions regarding ground truth often fail to hold, essentially to provide a principled way to synthesize granular observations into a cohesive measure of cultural intelligence. We outline below these challenges in more detail and discuss key strategic considerations for tackling them. Our goal is not to offer a singular, universal solution, but to provide a roadmap to measure various cultural intelligence indicators.
@@ 4.1 The Cultural Knowledge Frontier
Establishing a reliable, broad-coverage, and fine-grained cultural knowledge repository is foundational to a robust cultural intelligence evaluation suite. Such repositories of high fidelity cultural knowledge have already been used to serve as an evaluative foundation for AI generations; e.g., knowledge about cuisine and festivals [55], stereotypes [41], cultural norms [63, 69], to name a few. We note that this is essentially an endeavor of populating the “vocabulary of culture” (Section 2), and cultural knowledge here encompasses artifacts across both material and non-material facets. By structuring culture as a dynamic knowledge base consisting of rudimentary building blocks or artifacts, rather than a fixed set of labels on model inputs and outputs, we posit the creation of evaluation probes that are both extensible and deeply grounded. Instead of pursuing a “once and done” test of cultural appropriateness, this repository can be leveraged to generate controllable probes that precisely target a model’s understanding of specific cultural nuances.
Sourcing Cultural Knowledge: Populating the Vocabulary
@@
Populating the vocabulary of culture, even within any single topic or facet, is non-trivial. Current efforts to source global scale cultural knowledge frequently come at the expense of depth, while granular investigations into specific traditions are often impossible to scale. Consequently, foundational knowledge-gathering remains geographically siloed within a few high-resource cultural contexts, while grassroots efforts to capture under-represented cultures remain fragmented. Bridging this gap requires a unified approach to sourcing diverse cultural knowledge, particularly the “long-tail” aspects that are often invisible to digital crawlers and standard AI datasets. To resolve this friction, we envision a repeatable, modular, and multi-pronged data collection approach designed to balance the efficiency of automated methods with the indispensable authenticity of human lived experience. Such efforts should adopt a layered structure, allowing for a progression from efficient automated methods to increasingly complex and nuanced human curation.
@@
  • First, leveraging large-scale knowledge bases like Wikidata offers a path toward curating a reliable cultural repository with broad initial coverage (e.g., [43]), especially for stable topics of the Vocabulary of Culture such as Landmarks or Festivals. However, these sources often harbor significant data gaps, particularly regarding underrepresented cultures and/or in digitally underserved languages. The reliability of the information present in such existing knowledge bases can also vary substantially across different cultural contexts and facets [18].
@@
  • Some of these data gaps could be addressed through targeted LLM probing (e.g., [41]), often utilizing zero- and few-shot prompting that elicits the latent sociocultural knowledge captured within these models. This knowledge may have been missing from formal databases or have regional variants such as stereotypes or cuisine. However, this approach introduces its own set of sociotechnical risks: LLM-generated knowledge is frequently compromised by hallucinations, stereotypical biases, and spurious correlations that can distort cultural realities. Furthermore, using LLM-derived data to evaluate other LLMs creates a dangerous circularity, where models are essentially benchmarked against their own internal biases and errors. Consequently, rigorous manual validation by human experts with situated sociocultural knowledge [41] is not merely an enhancement, but a critical safeguard.
@@
  • While knowledge bases and LLM probing provide breadth, they both are limited by the “digital shadow” of the internet: they can only reflect what has already been digitized. Direct community engagement addresses this gap by capturing the nuanced, “offline” cultural knowledge that remains absent from web-scale datasets [70, 40]. This approach essentially surfaces the intra-regional long-tail: the hyper-local customs, rituals, dialects, and social norms that exist within local communities. Since this knowledge is underrepresented in model training data, it serves as a high-signal subset that reveals critical points of failure where models rely on shallow stereotypes or Western-centric defaults. However, moving into community spaces necessitates a rigorous and reciprocal ethical framework [12] that prioritizes community agency, active consent, and epistemic humility to ensure cultural knowledge is gathered non-extractively and non-exploitatively.
The resultant foundational resource created through this framework of strategies used in tandem can not only ground current evaluations but can serve as a controllable engine to develop the next generation of culturally intelligent AI probes [32, 41, 84].
Structuring Cultural Knowledge: From Rigid to Fluid Schemas
@@
Unlike standard knowledge bases optimized for objective ‘triplets’ (e.g., Paris is the capital of France), cultural knowledge is inherently more nuanced, pluralistic, and fluid. Technically, this introduces the “ground truth” problem: any specific artifact holds different meanings and importance depending on the observer’s intersectional identity or location. For instance, stereotypes about certain identities vary from place to place [41]; stability of consensus on cultural norms changes from time to time [22]. Representing these nuances may require moving away from rigid. monolithic taxonomies toward a hybrid architecture that balances the structural integrity of graph databases with the nuanced associations of vector-based representations, enabling us to model aspects such as the provenance and the temporal stability. Navigating this tension between algorithmic rigidity and cultural fluidity remains a crucial research pathway on this topic.
@@ 4.2 Probing for Cultural Intelligence Indicators
Another crucial consideration for ensuring prompt diversity is reflecting a wide range of user needs and incorporating various linguistic and structural variations  [27, 102] The set of prompts should reflect not only a diversity of language families  [75], but also the languages that are relevant in different geographic regions (e.g. English and Akan in Ghana in addition to other commonly used languages such as Ewe, Dagbani, and Hausa). While variations in sentence structure operate along the syntagmatic axis, prompt perturbations should also see variation along the paradigmatic axis, where relations between words are those that can be expressed in a taxonomy or substituted for each other. For example, demographics such as gender, age and education level result in lexical norms that differ between groups [87, 90]. When probing cultural information, diversifying prompts along all these axes can ensure that different styles of writing and interaction are represented as part of the probing method and the measurements are robust to such variations [4, 99].
@@ 4.3 Quantifying Cultural Capabilities: Metrics and Aggregation
The question of devising targeted prompts for cultural intelligence goes hand in hand with how to address the practical challenges of scoring them both effectively and efficiently. This transition involves navigating critical implementation trade-offs regarding robustness and cost, the selection of diverse scoring approaches, and the necessity of metric innovation where generic approaches fall short. Furthermore, we must consider how these varied measurements can be systematically aggregated across capabilities to provide a holistic and interpretable assessment of a model’s cultural intelligence.
Scoring Methodologies: Human, Objective, and Automated
@@
Because cultural intelligence encompasses diverse capabilities with a high degree of subjectivity in ground truth, a “one-size-fits-all” scoring method is rarely sufficient [6]. Instead, a strategic mix of approaches is required to achieve both nuance and scale. Human ratings remain the gold standard for perception-based indicators, as they capture the subjective social judgments necessary for evaluating qualities like cultural fluency, appropriateness, and coherence. However, metrics to effectively incorporate the inherent subjectivity in such socio-culturally shaped concepts is an active area of research [35, 60, 56]. In contrast, direct objective measures can be derived straight from model outputs, such as verifying factual correctness against localized knowledge repositories or quantifying intrinsic properties like response entropy [44]. Representing a scalable compromise, auto-raters or LLM-as-a-judge, i.e., specialized AI models trained to mimic human judgment, offer increasingly close approximations for human nuance at a fraction of the cost [96, 50]. These raters also facilitate objective measures by performing component tasks, such as entity extraction or semantic alignment with cultural databases. These methodological choices entail significant trade-offs. Foremost among these is the circularity inherent in using auto-raters, i.e., employing AI systems to evaluate the capabilities of other AI models. It thus requires rigorous safeguards to prevent measurement leakage, data contamination, and the perpetuation of model-centric biases, ensuring that the evaluation is robust against these factors. Furthermore, a strategic choice must be made regarding the foundational scoring methodology, deciding when to prioritize human nuance over automated scale.
The Need for Metric Innovation
@@
Evidently, evaluating cultural intelligence requires more than the application of general-purpose NLP benchmarks; it necessitates a fundamental innovation in how we define and measure “correctness” [39]. For instance, standard metrics used in tasks like question-answering or summarization, e.g., BLEU, ROUGE, or BERTScore [23], are designed to reward semantic similarity and lexical overlap. However, they may not port over to cultural evaluation readily because they are not sensitive to the nuances of cultural context — a model’s response might achieve a near-perfect similarity score against a reference answer yet fail critically due to a subtle misuse of honorifics, an inappropriate register, or the omission of a localized social convention. Conversely, two responses may be semantically distant but equally valid depending on the specific cultural sub-group or dialect being addressed. This discrepancy highlights the need for metrics that can distinguish between factual accuracy and pragmatic alignment. For example, a model might correctly translate a request for help but use a tone that is considered imperious or face-threatening in a high-context culture. Furthermore, innovation is required to capture the plurality of valid responses. Unlike standard question answering evaluations, where there is often a singular ground truth, cultural intelligence involves navigating ambiguity and multiple “correct” ways of being (e.g., [60]). New metrics should move away from rigid reference-based matching toward probabilistic or evaluative frameworks that account for cultural variation. This includes developing specialized metrics that prioritize the preservation of social meaning over lexical precision.
@@
Synthesizing these diverse indicators into a holistic assessment is a non-trivial task. Since a single number, or even a static array of numbers, often fails to capture the multi-faceted nature of cultural intelligence, we argue that aggregation must be goal-oriented, with different combinations of evaluations providing distinct perspectives based on the specific use case. The capability-indicator-measurement hierarchy of our framework, grounded in the work of Adcock and Collier [1], provides a systematic basis for aggregating evaluation results. By aligning specific indicators with their corresponding higher-level constructs, the framework ensures that aggregations are both mathematically coherent and conceptually valid The most straightforward approach involves intra-capability aggregation, where performance is averaged across cultures or tasks, potentially weighted by user-defined priorities or the empirical validity of specific tasks. Aggregating across different capabilities is more challenging and requires converting diverse task outputs into commensurable forms. For instance, rather than comparing raw scores, researchers might measure the expected errors per $N$ queries or the “distance to first error” across all capabilities to create a unified baseline. Finally, for head-to-head model analysis and grading, a granular query-level comparison tracking the frequency with which one model outperforms, equals, or underperforms another across all evaluations—provides the clearest picture of relative utility in cross-cultural contexts.
@@ 5 Discussion
Quantifying a construct as multifaceted as culture necessitates a strategic reduction of its complexity. While we deliberately attempted a more elemental and tractable approach by systematizing the concept of cultural intelligence to enable its measurement in the context of AI systems, such an endeavor is inherently subject to limitations and caveats. We start by discussing some of the key differentiators of our approach, followed by the limitations and ethical considerations inherent in this work.
@@ 5.1 Key Differentiators of Our Approach
Our framework for measuring cultural intelligence is anchored in a structured approach that centers on a fine-grained operationalization of culture and a capability-indicator-measurement conceptualization of cultural intelligence. This architectural clarity enables several key advantages: it allows for the seamless expansion of cultural scope across time and contexts, provides granular control over evaluation objectives, and yields actionable insights from discovered model losses.
@@ 5.2 Limitations
Inherent incompleteness of the cultural knowledge repository: Our reliance on a cultural knowledge repository is a double-edged sword. The reliability and fidelity of our measurement framework are thus directly dependent on the accuracy, coverage, and completeness of this underlying data [66]. While we have outlined methods to continually expand the repository, it will never be complete. Culture is infinite and dynamic; thus, any assessment grounded in a finite dataset is necessarily a sample of cultural intelligence, not an exhaustive evaluation. This limitation demands humility in interpretation and requires thoughtful investment in data maintenance and expansion [71].
Biases in data: Societal biases often creep into various stages of the data sourcing pipeline that will cause the repository to reflect some of these biases in coverage, which will then carry over to biases in the reliability of our measures. For instance, existing knowledge bases have been shown to have numerous biases in terms of their coverage [78], so if data is sourced from such existing knowledge bases, it is important to make sure they are extended through other complementary sources such as language models and/or community engagements. To address this limitation, it is important that we also continue to analyze for representational biases. Furthermore, any sampling from the data repository for evaluation datasets, could take into account balancing the prompts across dimensions that the use case calls for.

### v2

@@ 3 Operationalization: Indicators of Cultural Capabilities
Broadly, these indicators fall into two categories that balance objective verification with subjective human judgment: The first type is knowledge-based indicators, that measures the correctness and coverage of model responses against verifiable external truth sources. The goal here is to check the model’s factual recall and the fidelity of its representation of established, explicit cultural information. For example, a knowledge-based assessment might query a model on the recognized meaning of a specific national holiday, the ingredients of a regional dish, the historical significance of a landmark, detect the cultural rituals depicted in an image. Since these assessments seek answers against a known “ground truth,” they often provide an objective, scalable measure of the model’s acquired knowledge. (See Section 4 for the discussion on where scoring is a challenge even when there exist objective answers.) The second, more nuanced type is the perception-based assessment that recognizes the aspects of cultural intelligence that resides in pragmatics and context, which cannot usually be checked against a simple database. It relies on evaluating model responses using fluid, often subjective rubrics that capture perceptions of quality and fluency beyond mere factuality. These indicators move the questions from the more objective “is the answer right?” to the more subjective “is the answer appropriate?”, “is the answer fluent?” etc. Evaluations to measure such indicators usually require human ratings (or LLM-as-judge models trained or tuned on human ratings), and are crucially dependent on the representational diversity of the human rater pool itself [65, 66, 67]. Without diverse raters, a model might be rated as “fluent” simply because it aligns with the dominant culture’s expectations, while failing to resonate with the target community.
@@
By strategically combining these knowledge-based and perception-based indicators, one can construct a robust, multi-layered “thick evaluation” [68], that moves beyond the “thin” view of mere factual accuracy or trivia questions, which may confirm a model recognizes a cultural artifact, to probe the deeper, sociocultural resonance of its outputs. This comprehensive framework ensures that the resulting measure reflects not only what a model knows, but also how effectively and safely it can act across diverse cultural contexts.
@@
Translating abstract capabilities into specific behavioral or output-based markers yields us a few key advantages. First, it offers the flexibility for developers to weigh different indicators differently depending on the overarching product goals. For instance, the global diversity indicator may be de-prioritized in the context of a model that is aimed to be launched within a specific country, while stereotype avoidance may be an indicator that is chosen based on the normative objectives of the developer. Second, it offers a reproducible methodology to assess cultural intelligence in an extensible manner. As AI models gain new modalities and as cultural norms evolve over time, our framework allows for the continuous integration of new indicators without requiring a total overhaul of the underlying capability definitions, enabling our evaluation methods to remain relevant, rigorous, and responsive over time.
@@ 4 Measurement
We now move to the practical considerations of measuring these indicators at scale, navigating the complex array of technical and sociotechnical challenges. For instance, measuring Factual Accuracy, an indicator of Epistemic Fidelity, effectively requires us to possess: first the ground-truth knowledge the model is being tested for; second, a robust a set of probes (or prompts) that necessitate the model to use or surface that fact; and finally, a metric that can assess how well the model’s response matches the ground truth. Consequently, moving from the theoretical framework to a functional measurement requires navigating the friction between the fluid nature of culture and the relatively more rigid requirements of algorithmic benchmarking. We outline key strategic considerations for tackling these challenges. Our goal is not to offer a singular, universal solution, but to provide a roadmap to measure various cultural intelligence indicators.
@@ 4.1 The Cultural Knowledge Frontier
Establishing a reliable, broad-coverage, and fine-grained cultural knowledge repository is foundational to a robust cultural intelligence evaluation suite. By structuring culture as a dynamic knowledge base consisting of rudimentary building blocks or artifacts, rather than a fixed set of labels on model inputs and outputs, we posit the creation of evaluation probes that are both extensible and deeply grounded.
Sourcing Cultural Knowledge
@@
Populating the vocabulary of culture is non-trivial. Current efforts to source global scale cultural knowledge frequently come at the expense of depth, while granular investigations into specific traditions are often impossible to scale. To tackle this, we envision a repeatable, modular, and multi-pronged data collection approach designed to balance the efficiency of automated methods with the indispensable authenticity of human lived experience.
@@
  • First, leveraging large-scale knowledge bases like Wikidata offers a path toward curating a reliable cultural repository with broad initial coverage (e.g., [21]). However, these sources often harbor significant data gaps in coverage and quality/reliability, particularly regarding underrepresented cultures and/or in digitally underserved languages [69].
@@
  • Targeted LLM probing (e.g., [60]) could address some of these data gaps, often utilizing zero- and few-shot prompting that elicits the latent sociocultural knowledge captured within these models. This knowledge may be missing from formal databases. However, this approach introduces its own set of sociotechnical risks such as hallucinations, stereotypical biases, and spurious correlations. Rigorous manual validation by human experts with situated sociocultural knowledge [60] is hence not merely an enhancement, but a critical safeguard.
@@
  • While knowledge bases and LLM probing provide breadth, they both are limited as they can only reflect what has already been digitized. Direct community engagement addresses this gap by capturing the nuanced, “offline” cultural knowledge that remains absent from web-scale datasets [40, 61]. This approach can surface long-tail facets such as hyper-local customs, rituals, dialects, and social norms. Since this knowledge is underrepresented in model training data, it serves as a high-signal subset that reveals critical points of failure where models rely on shallow stereotypes or Western-centric defaults.
@@
Unlike standard knowledge bases optimized for objective ‘triplets’ (e.g., Paris is the capital of France), cultural knowledge is inherently more nuanced, pluralistic, and fluid. Technically, this introduces the “ground truth” problem: any specific artifact holds different meanings and importance depending on the observer’s intersectional identity or location. Representing these nuances may require moving away from rigid. monolithic taxonomies toward a hybrid architecture that balances the structural integrity of graph databases with the nuanced associations of vector-based representations, enabling us to model aspects such as the provenance and the temporal stability.
@@ 4.2 Probing for Cultural Intelligence Indicators
Another crucial consideration for ensuring prompt diversity is reflecting a wide range of user needs and incorporating various linguistic and structural variations  [72, 73] The set of prompts should reflect not only a diversity of language families  [74], but also the languages that are relevant in different geographic regions. While variations in sentence structure operate along the syntagmatic axis, prompt perturbations should also see variation along the paradigmatic axis, where relations between words are those that can be expressed in a taxonomy or substituted for each other. When probing cultural information, diversifying prompts along all these axes can ensure that different styles of writing and interaction are represented as part of the probing method and the measurements are robust to such variations [75, 76].
@@ 4.3 Quantifying Cultural Capabilities: Metrics and Aggregation
The question of devising targeted prompts for cultural intelligence goes hand in hand with the practical challenges of scoring them both effectively and efficiently.
@@
Because cultural intelligence encompasses diverse capabilities with a high degree of subjectivity in ground truth [77], a strategic mix of approaches is required to achieve both nuance and scale. While human ratings remain the gold standard for perception-based indicators that require subjective social judgments, metrics to effectively incorporate the inherent subjectivity in an AI evaluation context is an active area of research [78, 79, 80]. In contrast, direct objective measures can be derived straight from model outputs, such as verifying factual correctness against localized knowledge repositories or quantifying intrinsic properties like response entropy [81]. Representing a scalable compromise, LLM-as-a-judge approaches (i.e., specialized AI models trained to mimic human judgment) offer increasingly close approximations for human nuance at a fraction of the cost [82, 83]. These choices entail significant trade-offs in deciding between human nuance vs. scale. Furthermore, the circularity of employing AI systems to evaluate the capabilities of other AI models requires rigorous safeguards to prevent measurement bias and data contamination.
@@
Evidently, evaluating cultural intelligence requires more than the application of general-purpose NLP benchmarks; it necessitates bridging gaps in how we define and measure “correctness” [84]. Standard metrics used in tasks like question-answering or summarization, e.g., BLEU, ROUGE, or BERTScore [85], are designed to reward semantic similarity and lexical overlap. However, they may not port over to cultural evaluation readily because they are not sensitive to the nuances of cultural context; e.g., a model’s response might achieve a near-perfect similarity score against a reference answer yet fail critically due to a subtle misuse of honorifics, an inappropriate register.This discrepancy highlights the need for metrics that can distinguish between factual accuracy and pragmatic alignment. Furthermore, innovation is required to capture the plurality of valid responses. Unlike standard question answering evaluations, where there is often a singular ground truth, cultural intelligence involves navigating ambiguity and multiple “correct” answers.
@@
Since a single number, or even a static array of numbers, often fails to holistically capture the multi-faceted nature of cultural intelligence, we argue that aggregation must be goal-oriented, with different combinations of evaluations providing distinct perspectives based on specific use cases. The capability-indicator-measurement hierarchy of our framework provides a systematic basis for aggregating evaluation results. By aligning specific indicators with their corresponding higher-level constructs, the framework ensures that aggregations are both mathematically coherent and conceptually valid. Aggregating across different capabilities is challenging and requires converting diverse task outputs into commensurable forms. For instance, rather than comparing raw scores, researchers might measure the expected errors per $N$ queries or the “distance to first error” across all capabilities to create a unified baseline. Finally, for head-to-head model grading, a granular query-level comparison tracking the frequency with which one model outperformsanother provides the clearest picture of relative utility across cultural contexts.
@@ 5.1 Key Differentiators of Our Approach
Our framework is anchored in a structured approach that centers on a fine-grained operationalization of culture and a capability-indicator-measurement conceptualization of cultural intelligence. This architectural clarity enables several key advantages: extensibility across time and contexts, granular control over evaluation objectives, and actionable insights from discovered model losses.
@@ 5.2 Limitations
Our reliance on a cultural knowledge repository is a double-edged sword. The reliability and fidelity of our measurement framework are thus directly dependent on the accuracy, coverage, and completeness of this underlying data [86]. While we have outlined methods to continually expand the repository, any assessment grounded in a finite dataset is necessarily a sample of cultural intelligence, not an exhaustive evaluation. This limitation demands humility in interpretation and requires thoughtful investment in data maintenance and expansion [87]. Societal biases often creep into various stages of the data sourcing pipeline that will cause the repository to reflect some of these biases in coverage, which will then carry over to biases in the reliability of our measures. For instance, existing knowledge bases have been shown to have numerous biases in terms of their coverage [88], so if data is sourced from such existing knowledge bases, it is important to make sure they are extended through other complementary sources such as language models and/or community engagements. To address this limitation, it is important that we also continue to analyze for representational biases. Furthermore, any sampling from the data repository for evaluation datasets, could take into account balancing the prompts across dimensions that the use case calls for.